Tai
is fencing a triangular garden plot. he wants5 posts along each side. What is the least numbewr of posts that Tai will need?

Answer by checkley79(3341)

The answer is 12 posts.
1 for each corner (3*1)=3
3 additional posts between the end posts (3*3)=9
3+9=12 ans.
Proof: draw a diagram.